OpenStudy (anonymous):

the length of a rectangle is 3 cm greater than its width. The perimeter is 24cm. What are the dimensions of the rectangle.

OpenStudy (bibby):

Perimeter = length + length + width + width length = width + 3 width = width

OpenStudy (nurali):

Perimeter = 2(W + L) 24 = 2(W + (W+3)) 12 = 2W + 3 2W = 9 Width = 4.5 cm Length = 4.5+3 = 7.5 cm

OpenStudy (anonymous):

thank you very much

OpenStudy (nurali):

My Pleasure.